The product of 3 times a number and twice the number decreased by 7 is 0. What is the number?
Answer by rfer(16322)

3x*2x-7=0
6x^2-7=0
x^2=7/6
x=sqrt 7/6
x=1.080
-----------
3*1.080*2*1.080-7=0
3.24*2.16-7=0
7-7=0